students grades in a computer science course are assessed according to their performance in assignment, a mid course exam and a final exam. the relative weightings of these are:

assignment 20%
mid course exam 25%
final exam 55%

a particular student marks of 52 out of 60 for the assignment, 48 out of 100 for the mid course exam and 155 out of 2000 for the final exam. based on the above weightings calculate a weight mark out of 100 for this student.
1
Expert's answer
2020-04-16T19:07:19-0400

Assignment marks = "\frac{52}{60}" (relative weighting 20%)


Mid course marks = "\frac{48}{100}" (relative weighting 25%)

Final exam marks = "\frac{155}{2000}" (relative weighting 55%)


All the marks are converting as out of 100,


Assignment marks = "\frac{52}{60}*100=\frac{260}{3}"


Mid course marks = "\frac{48}{100}*100=48"


Final exam marks = "\frac{155}{2000}*100=7.75"


Weight mark calculates by multiplying the weighting factors,


Weight mark = "\frac{260}{3}*\frac{20}{100}+48*\frac{25}{100}+7.75*\frac{55}{100}"

Weight mark= "\frac{52}{3}+12+\frac{341}{80}"

Weight mark= 33.60 %